OpenStudy (anonymous):

Elizabeth has 1 7/10 as may plants as Rosalie has in her garden. If Elizabeth has 51 plants, how many plants does Rosalie have in her garden. please tell me how you got the answer also

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

brainly.com/question/777323 says its 30 plants

OpenStudy (richie63):

30 f sure

OpenStudy (alivejeremy):

E = (17/10) R E = 51 so (17/10) R = 51 R = 510/17 = 30

OpenStudy (anonymous):

rosalie has unknown plants (x) elizerbeth has 17/10 . elizerbeth has 51 therefore see below diagram

OpenStudy (anonymous):